OpenStudy (anonymous):

simplify √2x - √ 8x + √18x

OpenStudy (anonymous):

\[\sqrt{8x}\rightarrow \sqrt{(4*2)x}\rightarrow 2\sqrt{2x}\] \[\sqrt{18x}\rightarrow \sqrt{(9*2)x}\rightarrow 3\sqrt{2x}\] \[√2x - 2√ 2x + 3√2x=2\sqrt{2x}\]

OpenStudy (anonymous):

ty so much